NGI DE CH Operations Center:Operations Meeting:16032012

From EGIWiki
Jump to: navigation, search

Operations Meeting Main

Introduction

  • Minutes of last meeting

Announcements

Feb:
VoOps:
Av/Re= 97%
UNI BONN only has 69%

BDII: Av/Re=99,3%
 
  • Monitoring
https://helpdesk.ngi-de.eu/index.php?mode=ticket_info&ticket_id=1969
https://helpdesk.ngi-de.eu/index.php?mode=ticket_info&ticket_id=1978
High rate of unknown Nagios tests. We also contacted sites MPI-K, UNI-Dresden, UNI-Karlsruhe, UNI-Siegen per email. Reason: May be 
connection problems
The NGI-DE Nagios was also down during our (KIT, FZKA-LCG2) downtime last week. We will recalculate the availability numbers for 
March.
  • Staged rollout/updates
UMD:
http://www.eu-emi.eu/emi-1-kebnekaise-updates/-/asset_publisher/Ir6q/content/update-14-16-03-2012
http://www.eu-emi.eu/emi-1-kebnekaise-updates/-/asset_publisher/Ir6q/content/update-13-17-02-2012
  • Survey
Usage and future maintenance of deployed software
Operations/Platform Deployment Survey
For the surveys we need your feedback untill 20st of March via Email

Round the sites

NGI-DE
  • BMRZ-FRANKFURT (Uni Frankfurt)
  • DESY-HH
  • DESY-ZN
  • FZJuelich
  • Goegrid
  • GSI
  • ITWM (Martin Braun)
 - 09.03 -> ITWM will not be able to attend the phone conference on 9/3/12. There is nothing special to report.
 - 16/3/12 ntr
  • KIT (GridKa, FZK-LCG2, Dimitri Nilsen, Tobias Koenig)
 - 9.03 -> Announcement: downtime 13.03 - 15.03, full site maintainance
 - 16/3/12 downtime succeeded with two hours delay. During the downtime one central router was replaced by a new model, also we did 
some dCache updates. disk firmware and BIOS upgrades, Tape TSM update, complete cluster was reinstalled, changes of the central 
power supply, some LFC and 3D-DB were migrated to Oracle version 11g
  • KIT (Uni Karlsruhe)
  • LRZ
  • MPI-K
  • MPPMU (Cesare Delle Fratte)
 - ntr
 - a few problems with CREAMS and ATLAS jobs. Problems are solved now.
  • RWTH Aachen
  • SCAI (Andre Gemuend)
 - Problems with SE: DPM daemon died. We will update EMI-DPM
 - ROD also filed two tickets (one concerning the DPM daemon problem and one concerning the related CREAM CE) instead of only one 
for the dpm daemon: 
  • Uni Bonn
  • Uni Dortmund
  • Uni Dresden (Ralph Mueller Pfeeferkorn)
 - 16/3/12 Last week we had a two days downtime. We updated CREAM, Apel to the EMI release, dCache was updated to version 1.9.12 
and dCache update included the upgrade from SL4 to SL5. Now all seems to be fine now 
 - Problems with EMI release: After the update sometimes the Nagios test fails, with error message "job could not be submitted" or 
something like that. Comment by KIT: We also tested the BDII in EMI in preproduction. It was not described in the documentation 
that you need 3GB RAM.
  • Uni Freiburg
  • Uni Mainz-Maigrid
  • Uni Siegen
  • Uni Wuppertal
SwiNG
  • CSCS
  • PSI
  • Switch

Note: please update your entry at https://wiki.egi.eu/wiki/NGI_DE:Sites if needed.

Status ROD

welcome LRZ Daniel Waldmann

09	 27.02	 04.03	 Team5, LRZ	
10	 05.03	 11.03	 Team6, CSCS/NGI_CH	
11	 12.03	 18.03	 Team1, DESY	
12	 19.03	 25.03	 Team2, FhG
  • Nagios<-->Dashboard issue
(Nagios send information/notification to the dashboard and the information/notification is not displayed correctly at the dashboard. 
In consequence the tickets hang in the dashboard longer than the problem persists)
 https://helpdesk.ngi-de.eu/?mode=ticket_info&ticket_id=2039
 https://helpdesk.ngi-de.eu/?mode=ticket_info&ticket_id=2038

AOB

 - Reminder: Operation workshop in April: We hope many sites participate. You can register now. See announcement section above.
 - short Telco meeting before the workshop at 13. April

If you have additional topics to be discussed during the meeting, please submit them in advance via our email list email list.